OpenStudy (anonymous):

which means the sum of w and 3.4 is greater than or equal to 10.5

OpenStudy (saifoo.khan):

\[w + 3.4 \ge 10.5\]

OpenStudy (anonymous):

\[w+3.4\ge10.5\]

OpenStudy (saifoo.khan):

\[w \ge 10.5 - 3.4 \]

OpenStudy (saifoo.khan):

\[w \ge 7.1\]

OpenStudy (anonymous):

\[w+3.4\ge10.5\]

OpenStudy (anonymous):

\[w \ge7.1\]